# Deeper, Sharper, Faster: Application of Efficient Transformer to Galaxy Image Restoration

### v1

@@ Abstract
facilities: JWST (NIRCam), HST (ACS)software: numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), astropy (Astropy Collaboration et al., 2013, 2018), photutils (Bradley et al., 2023), SExtractor (Bertin & Arnouts, 1996), GalSim (Rowe et al., 2015)
@@ 1 Introduction
Deep learning techniques offer advantages over traditional methods, such as Fourier deconvolution and Bayesian approaches, by providing greater flexibility and adaptability to diverse datasets. The integration of CNNs with other deep learning architectures, such as generative adversarial networks  (GANs; Goodfellow et al., 2014) and recurrent neural networks  (RNNs; Williams & Zipser, 1989), has further expanded the capabilities of image enhancement  (e.g., Ledig et al., 2016; Schawinski et al., 2017; Liu et al., 2021; Tripathi et al., 2018; Alsaiari et al., 2019; Rajeev et al., 2019; Wang et al., 2020; Tran et al., 2021; Kalele, 2023). These hybrid approaches enable the generation of highly realistic and detailed images, pushing the boundaries of what is achievable with traditional methods alone. One of the outstanding limitations of the CNN-based model is its restricted receptive field. That is, long-range pixel correlations may not effectively be captured by the model. Another critical drawback is its static weights, which prevent effectively adapting to input content during inference.
@@ 2.4 Implementation Details
We use a transfer learning approach, where a model trained on one dataset is reused for another related dataset. First, we train on the pre-training dataset (simplified galaxy images) for 150,000 iterations, followed by an additional 150,000 iterations on the finetuning dataset (realistic galaxy images). The batch size remains fixed at 64. Additionally, to compare and analyze the performance of training on individual datasets, we also conduct training separately solely based on either the pre-training or finetuning datasets. Our inference model is publically available  [^1] https://github.com/JOYONGSIK/GalaxyRestoration.
@@ 3 Data
The GT image of the finetuning dataset is obtained by collecting galaxy images from the JWST archive. Following the same procedure in the pre-training dataset generation, we create their LQ counterparts. Below we describe the HST, GalSim, and JWST datasets in detail.
@@ 3.1 HST Dataset
The first and second criteria eliminate excessively large or small sources. The third condition removes excessively faint sources. The fourth condition excludes the cases where similarly bright galaxies are present within the postage stamp image. The last condition is necessary because the HST automatic cosmic ray detection/removal pipeline often fails near the bright peaks of stars or galaxies. In addition to these criteria, we explicitly identified and removed stars using their distinct locus in the size-magnitude diagram. We obtain a total of 66,092 postage-stamp images. Figure 3 shows the distribution of the HST postage-stamp images prior to the $w=128$ cut.
@@ 3.3 JWST Dataset
We standardize the postage-stamp image size to $64\times 64$ pixels. A smaller image is expanded by padding pixels to its four edges in such a way that both the noise and background levels match the original image. A larger image is trimmed to $64\times 64$ pixels. We discarded the galaxy images whose profiles were prematurely truncated by the trimming procedure. We collected 113,485 postage-stamp galaxy images from the JWST NIRCam F115W, F200W, F277W, and F [^4] W images from various fields publicly available in August 2023 the Mikulski Archive for Space Telescopes (MAST) [^4] https://archive.stsci.edu/. The dataset serves as the GT component in our finetuning dataset.
The LQ counterpart is generated by convolving the JWST GT galaxy images with the HST PSF and adding Gaussian noises. Although the pixel scale is different from the HST image, we scaled the HST PSF size, assuming the $0.05\arcsec$ pixel scale. This is because, eventually, we desire to restore the HST images whose pixel scale is $0.05\arcsec$. The applied noise distribution follows that of the HST dataset. The ratios among the training, validation, and test datasets are 8:1:1.
@@ 4 JWST Test Dataset Results
In this section, we present the results when our deep learning model is applied to the JWST test dataset. We first define the two evaluation metrics: peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM), describe our visual inspection, and compare morphological and photometric properties.
@@ 4.1 PSNR and SSIM
where $\mu_{x}$ ($\mu_{y}$) and $\sigma_{x}^{2}$ ($\sigma_{y}^{2}$) are the mean and variance, respectively, of the $x$ ($y$) image, and $\sigma_{xy}^{2}$ is the covariance between the $x$ and $y$ images. $C_{1}$ and $C_{2}$ are introduced to prevent the equation from diverging when the denominator is small, and the exact values depend on the dynamic range of the target images. SSIM is a perceptual metric that evaluates the structural similarity between two images and considers luminance, contrast, and structure. The value ranges from -1 to 1 with 1 being the perfect similarity. In Table 1, we compare the PSNR and SSIM values between our fiducial model (transfer learning) and experimental (non-transfer learning) models. The transfer learning model provides the highest values in both PSNR and SSIM.
@@
Note. — The model obtained with transfer learning (bold) shows improvements in both PSNR and SSIM.
@@ 5 APPLICATION TO REAL HST IMAGES
Figure 8: Application to HST images. We apply our restoration model discussed in §2 without any modification to HST images. (a) Comparison of single-epoch RS images with multi-epoch high-quality images. Here, we use deep multi-epoch images as a proxy for GT. The single-epoch LQ images were restored using our model. Note that since the model is designed to perform deconvolution to match the JWST PSF, the resolution of the RS images is higher than that of the deep multi-epoch HST images. Also, the noise levels in the RS images are lower than in the multi-epoch data. (b) Comparison of multi-epoch RS images with deep JWST images. Here, we use deep JWST images as a proxy for GT. Since the JWST/NIRCam images are Near IR (F115W) images while the HST/ACS images are optical (F606W) images, only limited comparison is possible. Nevertheless, both the reduction in noise and improvement in resolution make the RS images look more similar to the JWST images than the corresponding LQ images.
@@
5.1 Restoration of Single-epoch Images and Comparison with Multi-epoch Images
We utilize the datasets from the HUDF and GOODS-N&\&&S observed by the HST ACS. The single-epoch low-quality (SG-LQ) images are retrieved from the MAST archive. We aim to restore the FLT images not their “drizzled” (DRZ)images because “drizzling” induces inter-pixel noise correlations. When a postage-stamp image is contaminated by cosmic rays, we remove them with the lacosmic script [^6] https://lacosmic.readthedocs.io/en/stable. Since the multi-epoch high-quality (MT-HQ) mosaic images are corrected for geometric distortions unlike the FLT images, we transformed them back to the distorted, CCD coordinate system to enable a fair comparison with the FLT images.
In Figure 8a, we display four restoration examples. Similar to the results observed in the experiment with the JWST test dataset, the restoration qualities are similarly excellent here. The low-surface brightness features, which cannot be identified in the LQ images, are well-restored in the RS images. One notable difference from the results in the JWST test dataset is the PSF size. Since our deep learning model is trained to restore the JWST resolution, the PSF size of the RS images are smaller than the HQ images. Another important difference is on average the higher noise level in our single-epoch LQ images.
@@ 5.1 Restoration of Single-epoch Images and Comparison with Multi-epoch Images
5.2 Restoration of Multi-epoch HST Images and Comparison with Multi-epoch JWST Images
@@
Figure 8b shows four examples. Despite the difference in filter, the RS images look remarkably similar to the JWST images. In particular, when the noise level is relatively low (i.e., top two panels), the restoration in both resolution and low-surface brightness feature is excellent. When the noise level is relatively high, the RS images still resemble the JWST images more closely than their input HST images. We suspect that the correlated noise in the input multi-epoch HST image might have non-negligible impact on the result in this noise regime.
@@ 6.3 Artifacts Due to Pixel Correlation
Figure 13 display some examples of these artifacts. The LQ images here are sampled from multi-epoch drizzled images. The presence of correlated noise is apparent even from visual inspection. The RS images show that the correlated noise creates some low-surface brightness artifacts in the galaxy outskirts, which however are absent in the JWST images. Addressing this issue could involve strategies such as employing a different drizzling kernel for image stacking or leveraging more advanced deep learning algorithms. Exploring these solutions will be a key focus of our future work.
Figure 13: Application to HST galaxies sampled from multi-epoch drizzled images. Since the drizzling algorithm induces substantial noise correlation, the input data have different noise properties from our training datasets. The RS images show that the correlated noise creates some low-surface brightness artifacts in the outskirts.
@@ 7 Conclusions
With the test dataset, we find that the restored images show significantly enhanced correlations with the GT images than their original LQ versions, reducing the scatters of isophotal photometry, Sersic index, and half-light radius by factors of 4.4, 3.6, and 4.7, respectively, with Pearson correlation coefficients approaching unity. We also visually confirm that the restored images are superior in terms of resolution and noise level. When we applied our model to real low-exposure HST images, the restored images also show significantly improved correlations with their multi-exposure versions, although the absence of their real GT images limits our interpretations.

### v2

@@ Abstract
facilities: JWST (NIRCam), HST (ACS)software: numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), astropy (Astropy Collaboration et al., 2013, 2018), photutils (Bradley et al., 2023), SExtractor (Bertin & Arnouts, 1996), GalSim (Rowe et al., 2015), lacosmic (Bradley, 2023), Drizzlepac(Hoffmann et al., 2021)
@@ 1 Introduction
Deep learning techniques offer advantages over traditional methods, such as Fourier deconvolution and Bayesian approaches, by providing greater flexibility and adaptability to diverse datasets. The integration of CNNs with other deep learning architectures, such as generative adversarial networks  (GANs; Goodfellow et al., 2014) and recurrent neural networks  (RNNs; Williams & Zipser, 1989), has further expanded the capabilities of image enhancement  (e.g., Ledig et al., 2016; Schawinski et al., 2017; Liu et al., 2021; Tripathi et al., 2018; Alsaiari et al., 2019; Rajeev et al., 2019; Wang et al., 2020; Tran et al., 2021; Kalele, 2023). These hybrid approaches enable the generation of highly realistic and detailed images, pushing the boundaries of what is achievable with traditional methods alone. One of the outstanding limitations of the CNN-based model is its restricted receptive field. That is, long-range pixel correlations may not effectively be captured by the model. Another critical drawback is its limited adaptability to input content during inference.
@@ 2.4 Implementation Details
Every input image is min-max normalized. This normalization scheme is needed because astronomical images span a wide dynamic range. Without the proper normalization, the training will be biased toward restoring the brightest galaxy images. Also, we find that this standardization helps to train our model more efficiently. Finally, the min-max normalization minimizes the effects of field-dependent calibration systematics.
We use a transfer learning approach, where a model trained on one dataset is reused for another related dataset. First, we train on the pre-training dataset (simplified galaxy images) for 150,000 iterations, followed by an additional 150,000 iterations on the finetuning dataset (realistic galaxy images). The batch size remains fixed at 64. Additionally, to compare and analyze the performance of training on individual datasets, we also conduct training separately solely based on either the pre-training or finetuning datasets. Our inference model is publically available  [^1] https://github.com/JOYONGSIK/GalaxyRestoration  [^2] Zenodo: https://doi.org/10.5281/zenodo.11378660 .
@@ 3 Data
The GT image of the finetuning dataset is obtained by collecting galaxy images from the JWST archive. Following the same procedure in the pre-training dataset generation, we create their LQ counterparts.
We provide detailed descriptions of the HST, GalSim, and JWST datasets in §3.1, §3.2, and §3.3, respectively. Below we briefly outline the purpose of each dataset.
  • HST dataset: To define the properties of galaxies for creating synthetic images and determine the appropriate noise level for generating low-quality images. This dataset is not used in training.
  • GalSim dataset: To pretrain the model on noise-free synthetic images of galaxies to learn core features.
  • JWST dataset: To finetune the model using real JWST galaxy images to learn diverse forms and characteristics of galaxies not covered by the GalSim dataset.
@@ 3.1 HST Dataset
The first and second criteria eliminate excessively large or small sources. The third condition removes excessively faint sources. The fourth condition excludes the cases where similarly bright galaxies are present within the postage stamp image. The last condition is necessary because the HST automatic cosmic ray detection/removal pipeline often fails near the bright peaks of stars or galaxies. In addition to these criteria, we explicitly identified and removed stars using their distinct locus in the size-magnitude diagram. These conditions define our domain of interest and ensure the purity of our dataset. On the other hand, the omission of these galaxies may restrict the domain in which the model can perform well. We obtain a total of 66,092 postage-stamp images. Figure 3 shows the distribution of the HST postage-stamp images prior to the $w=128$ cut.
@@ 3.3 JWST Dataset
We standardize the postage-stamp image size to $64\times 64$ pixels. A smaller image is expanded by padding pixels to its four edges in such a way that both the noise and background levels match the original image. A larger image is trimmed to $64\times 64$ pixels. We discarded the galaxy images whose profiles were prematurely truncated by the trimming procedure. We collected 113,485 postage-stamp galaxy images from the JWST NIRCam F115W, F200W, F277W, and F [^4] W images from various fields publicly available in August 2023 the Mikulski Archive for Space Telescopes (MAST) [^5] https://archive.stsci.edu/: 10.17909/crbp-3069. The dataset serves as the GT component in our finetuning dataset.
The LQ counterpart is generated by convolving the JWST GT galaxy images with the HST PSF and adding Gaussian noises. Although the pixel scale is different from the HST image, we scaled the HST PSF size, assuming the $0.05\arcsec$ pixel scale. This is because, eventually, we desire to restore the HST images whose pixel scale is $0.05\arcsec$. The applied noise distribution follows that of the HST dataset. The ratios among the training, validation, and test datasets are 8:1:1. The GalSim and JWST train set subsamples are publically available. [^6] https://drive.google.com/file/d/1dOemrQXFr2UxHgvUTUffLqIXlHeEkRbZ/view?usp=sharing
@@ 4 JWST Test Dataset Results
In this section, we present the results when our deep learning model is applied to the JWST test dataset. In this case, for each LQ image, there exists a corresponding GT image, which makes the comparison straightforward. However, readers are reminded that when we apply our model to real images (§5), where there are no matching GT images, we can only use GT proxies obtained also from real images for comparisons.
Below, we define the two evaluation metrics: peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) (§4.1), describe our visual inspection (§4.2), and compare morphological (§4.3) and photometric properties (§4.4).
@@ 4.1 PSNR and SSIM
where $\mu_{x}$ ($\mu_{y}$) and $\sigma_{x}^{2}$ ($\sigma_{y}^{2}$) are the mean and variance, respectively, of the $x$ ($y$) image, and $\sigma_{xy}^{2}$ is the covariance between the $x$ and $y$ images. $C_{1}$ and $C_{2}$ are introduced to prevent the equation from diverging when the denominator is small, and the exact values depend on the dynamic range of the target images. SSIM is a perceptual metric that evaluates the structural similarity between two images and considers luminance, contrast, and structure. The value ranges from -1 to 1 with 1 being the perfect similarity.
[…]
@@
no training 16.4488 0.1883
Note. — The model obtained with transfer learning (bold) shows improvements in both PSNR and SSIM. The last row presents the PSNR and SSIM measured directly between the GT and LQ datasets, displaying the initial quality of the datasets.
@@ 5 APPLICATION TO REAL HST IMAGES
Figure 8: Application to HST images. We apply our restoration model discussed in §2 without any modification to HST images. (a) Comparison of single-epoch RS images with multi-epoch high-quality images. Here, we use deep multi-epoch images as a proxy for GT. The single-epoch LQ images were enhanced using our model. Note that since the model is designed to perform deconvolution to match the JWST PSF, the resolution of the RS images is higher than that of the deep multi-epoch HST images. Also, the noise levels in the RS images are lower than in the multi-epoch data. (b) Comparison of multi-epoch RS images with deep JWST images. Here, we use deep JWST images as a proxy for GT. Since the JWST/NIRCam images are Near IR (F115W) images while the HST/ACS images are optical (F606W) images, only limited comparison is possible. Nevertheless, both the reduction in noise and improvement in resolution make the RS images look more similar to the JWST images than the corresponding LQ images.
@@
5.1 Enhancement of Single-epoch Images and Comparison with Multi-epoch Images
We utilize the datasets from the HUDF and GOODS-N&\&&S observed by the HST ACS. The single-epoch low-quality (SG-LQ) images are retrieved from the MAST archive. We aim to enhance the FLT images not their “drizzled” (DRZ)images because “drizzling” induces inter-pixel noise correlations. When a postage-stamp image is contaminated by cosmic rays, we remove them with the lacosmic script [^8] https://lacosmic.readthedocs.io/en/stable (van Dokkum, 2001; Bradley, 2023). Since the multi-epoch high-quality (MT-HQ) mosaic images are corrected for geometric distortions unlike the FLT images, we transformed them back to the distorted, CCD coordinate system to enable a fair comparison with the FLT images.
In Figure 8a, we display four enhancement examples. Similar to the results observed in the experiment with the JWST test dataset, the restoration qualities are similarly excellent here. The low-surface brightness features, which cannot be identified in the LQ images, are well-restored in the RS images. One notable difference from the results in the JWST test dataset is the PSF size. Since our deep learning model is trained to restore the JWST resolution, the PSF size of the RS images are smaller than the HQ images. Another important difference is on average the higher noise level in our single-epoch LQ images.
@@ 5.1 Enhancement of Single-epoch Images and Comparison with Multi-epoch Images
5.2 Enhancement of Multi-epoch HST Images and Comparison with Multi-epoch JWST Images
@@
Figure 8b shows four examples. Despite the difference in filter, the RS images look remarkably similar to the JWST images. In particular, when the noise level is relatively low (i.e., top two panels), the enhancement in both resolution and low-surface brightness feature is excellent. When the noise level is relatively high, the RS images still resemble the JWST images more closely than their input HST images. We suspect that the correlated noise in the input multi-epoch HST image might have non-negligible impact on the result in this noise regime.
@@ 6.3 Artifacts Due to Pixel Correlation
Figure 13 displays some examples of these artifacts. The LQ images here are sampled from multi-epoch images drizzled with square and Lanczos3 kernels using AstroDrizzle (Fruchter et al., 2010) in DrizzlePac (Hoffmann et al., 2021). With square-kernel drizzling, correlated noise is apparent even from visual inspection. The RS images show that the correlated noise creates some low-surface brightness artifacts in the galaxy outskirts, which are absent in the JWST images. These artifacts are mitigated when we use the Lanczos3 kernel, which helps reduce the pixel correlation in the mosaic-drizzled image. The background and edge of the galaxy in the RS images are clean and smooth. Addressing this issue could involve strategies such as employing a different drizzling kernel for image stacking or leveraging more advanced deep learning algorithms. Exploring these solutions will be a key focus of our future work.
Figure 13: Application to HST galaxies sampled from multi-epoch drizzled images. The first and third columns are the HST-LQ multi-epoch images drizzled with square and Lanczos3 kernels, respectively. Pixel values are significantly correlated for the LQ image produced with the square kernel. In their corresponding RS images, these correlations manifest as numerous artifacts, which are absent in the JWST-HQ images. The Lanczos3 kernel reduces pixel correlations in the LQ images and also minimizes artifacts in the corresponding RS images.
6.4 Discussion on Potential Hallucination Effects
A hallucination effect in generative models refers to the phenomenon where the model generates outputs that are not grounded in reality or the input data, but instead are fabricated or “hallucinated” by the model, leading to false interpretations. Since virtually all generative models are subject to hallucination effects in varying degrees, we discuss potential hallucination effects in our restoration model.
A detailed examination of Figure 8a reveals that the restored images of the single-epoch data in the third and fourth rows show some compact, point-like sources. Because these features do not exist in the corresponding much deeper multi-epoch data, they may be regarded as arising from hallucination effects. However, scrutiny of the corresponding LQ images indicates that the pixel levels of the regions where the hallucination effect is suspected are elevated compared to the neighboring regions. They may be outliers in noise fluctuations, uncorrected bad pixels, or residual cosmic rays that were not identified. We verified that these compact, point-like features do not appear in the RS images without their counterparts in the GT images when we restored images from the LQ images generated by degrading the JWST GT images (Figure 5). Therefore, we conclude that the artifacts in Figure 8a are grounded in the real features of the original LQ images rather than caused by the hallucination effect. Nevertheless, the discussion provides an important lesson: when we apply our model to real LQ images, we need to ensure that the input images do not possess glitches (e.g, bad pixels, cosmic rays, etc.) distinct from general noise.
@@ 7 Conclusions
With the test dataset, we find that the restored images show significantly enhanced correlations with the GT images than their original LQ versions, reducing the scatters of isophotal photometry, Sersic index, and half-light radius by factors of 4.4, 3.6, and 4.7, respectively, with Pearson correlation coefficients approaching unity. We also visually confirm that the restored images are superior in terms of resolution and noise level. When we applied our model to real low-exposure HST images, the enhanced images also show significantly improved correlations with their multi-exposure versions, although the absence of their real GT images limits our interpretations.